# Extreme events at the onset of epileptic-like chimeras in small-world networks of FitzHugh-Nagumo neurons

### v1

@@ Abstract
In this work, we investigate the dynamics of complex networks of FitzHugh-Nagumo excitable oscillators, focusing on the impact of coupling strength, network size, and randomness on their collective dynamics. Considering Watts-Strogatz small-world network connectivities, the system exhibits three distinct dynamical phases: chaotic, intermittent, and synchronized, with the intermittent phase displaying transient, epileptic-like chimera states. We analyse the transition to synchronisation by means of the master stability function, and show that peaks in the proportion of extreme events of synchronisation, which correlate with the behaviour of the largest Lyapunov exponent of the system, precede the transitions between the distinct dynamical regimes and mark the onset of epileptic-like chimera states. Our findings contribute to a broader understanding of synchronisation in excitable systems real neural networks and offer insights into the conditions that may lead to pathological epileptic-like states. Furthermore, we discus the potential use of extreme events to study real neural data.
@@ I Introduction
Synchronisation phenomena occur across diverse fields, ranging from physics and chemistry to biology, neuroscience, and even social systems [1, 2, 3, 4, 5]. In the brain, coordinated firing of distinct regions of the brain is crucial for effective information processing. However, when dysregulated, synchronized states can lead to pathological conditions such as epileptic seizures (see Fig. 1a), which disrupt normal brain function [6, 7]. Understanding the conditions that give rise to these coherent events, as well as the biological and physical mechanisms that regulate them, is therefore of considerable importance.
@@
The human brain is composed on billions of neurons, which are grouped and compartmentalized. These compartments establish an intricate functional connectome. Each region of the brain may be modelled as a single node, and interactions between regions as edges connecting nodes, such that the whole brain is reduced to a complex network [8, 9, 10, 11, 12] (see Fig. 1b), whose topology follows a small-world organization [13].
A paradigmatic model that captures the excitable and oscillatory regimes of neural systems is the FitzHugh-Nagumo (FHN) oscillator [14], which can be obtained as a qualitative reduction of the classical Hodgkin-Huxley model. This model exhibits a wide range of spatio-temporal phenomena, from extreme events on coupled neurons with time delay [15] and the emergence of Turing patterns in a reduced one dimensional coupled model [16], to chimera states in a non-locally coupled ring [17, 18, 19, 20, 21, 22, 23] and synchronisation phase transitions [20, 18, 24, 25]. When coupled through a complex network type of connectivity, chimera states have been related to epileptic seizures in the human brain. These states are characterized by a coexistence of coherent and incoherent phases, and the topology of the connectivity network may alter the duration of the coherent phases [6, 11, 12]. However, how the network properties and coupling strength control the emergence of these states and how the transition towards intermittent states can be predicted remain unclear.
@@
Figure 1: Networks of excitable systems. (a) Electroencephalogram (EEG) from a human brain exhibits synchronised regular activity between distinct regions of the brain, characteristic of an epileptic event (adapted from [26]). (b) Anatomical regions of the brain (left). Regions are interpreted as nodes, and the connectivity network (right) is inferred from the neural activity (adapted from [13]). (c) WS networks for different values of the rewiring probability ($p$), colour of the nodes indicate their degree ($k$). (d) Degree distribution for WS networks for different values of $p$ (increasing from left to right), and different network sizes (increasing from top to bottom).
In this work, we study the interaction between FHN oscillators coupled through a Watts-Strogatz (WS) small-world network, which is characterised by a high degree of clustering and small topological distance between nodes. There are several studies suggesting that real brain networks are not homogeneous, and may present small-world properties [9, 5, 8, 27]. WS networks are constructed starting from a ring lattice with $N$ nodes, with each node connected to its $\langle k\rangle$ nearest neighbours, these connections are rewired following a random selection rule with probability $p$ (see Fig. 1c). In this model, the probability of rewiring a connection $p$ accounts for the degree of randomness of the network: $p=0$ corresponds to a regular ring lattice, and $p=1$ corresponds to disordered network with small-world properties [5]. The degree distribution of these networks (see Fig. 1d) is affected by $p$ [28] and the network size $N$, and converges to a distribution peaked at $\langle k\rangle$ that decays exponentially away from it [28]. We note that this contrasts with power-law like decays of scale-free networks.
@@
Here, we show that the system of networked FHN oscillators has three distinct dynamical phases: a chaotic, an intermittent and a synchronised phase. The intermittent phase exhibits transient highly coherent, chimera states separated in time by chaotic behaviour, which were previously related to epileptic-like events [11]. We find that the network size and the level of randomness of its connectivity, as well as the coupling strength between oscillators control the emergence of epileptic-like chimera states. In particular, the network randomness controls the size of the region in the parameter space where epileptic-like events exist and are stable. By computing the master stability function (MSF) of the system, we characterise the transition from intermittency to synchronisation. Furthermore, by analysing the existence of extreme events in the time series of the Kuramoto global order parameter, we show that both the emergence of chimeras and the change in stability of the synchronised manifold are marked by an increase in the probability to observe short transients of anomalous high synchronisation, and we relate these to changes in the values of the largest Lyapunov exponent and dimension of the strange attractor.
@@
The paper is organised as follows: In Sec. I.1, we introduce the model, in Sec. II.1, we show that the frequency and duration of epileptic-like events depend on coupling strength, the number of oscillators, and network randomness. In Sec. II.2, we study the dynamical regimes in parameter space, as well as the stability of the synchronous solution. In Sec. II.3, we show that the transitions between dynamical regimes can be characterized by the likelihood of observing extreme synchronisation events. Finally, in Sec. III, we summarise and discuss our results, and their relevance to biological systems.
@@
Figure 2: Dynamical regimes in networks of FHN oscillators. Left panels show the time series of the Kuramoto order parameter $R(t)$ (top), and the temporal evolution of the phases $\theta_{j}(t)\in[-\pi,\pi]$ for each of the $j=1,...,N=50$ oscillators (bottom), right panels show the distribution of inter-spike intervals (ISI) for the $N=50$ oscillators, for coupling (a) $d=0.01$ (b) $d=0.05$ (c) $d=0.10$ (d) $d=0.22$. An epileptic-like event is highlighted in the left panels of (b).
I.1 The model
@@
The FitzHugh-Nagumo model, originally introduced by FitzHugh as a modification of the Van der Pol oscillator and later demonstrated by Nagumo to be analogous to an electrical circuit [29, 30], serves as a well-established framework for modelling excitable systems. The model comprises two variables: an activation variable, $x(t)$, which captures the rapid dynamics of excitation, and an inhibition variable, $y(t)$, representing the slower recovery process that relaxes the activity back to its basal level. In this study, we consider a network of $N$ coupled FHN oscillators, where the dynamics of each oscillator is governed by [31, 32, 22, 21]
$\left(\begin{array}[]{c}\varepsilon\dot{x}_{i}\\ \dot{y}_{i}\end{array}\right)=\left(\begin{array}[]{c}x_{i}-\frac{x_{i}^{3}}{3}-y_{i}\\ x_{i}+a\end{array}\right)+d\sum_{j=1}^{N}A^{p}_{ij}\bm{R}\left(\begin{array}[]{c}x_{j}-x_{i}\\ y_{j}-y_{i}\end{array}\right),$ (1)
@@
where $\dot{x}$ and $\dot{y}$ correspond to the time-derivatives of $x$ and y, respectively, and the time-scale separation $\varepsilon=0.05$ is fixed so that $x(t)$ has a fast dynamics compared to $y(t)$. The parameter $a$ controls a transition from an excitable ($|a|>1$) to an oscillatory ($|a|<1$) regime. As our interest is to study interactions of actively firing circuits, we set $a=0.5$. The sum term in Eq. (1) accounts for the inter-oscillator coupling, which is proportional to a uniform coupling strength $d$. The coupling matrix $\bm{R}$ is defined as
$\bm{R}=\left(\begin{array}[]{cc}\cos\theta&\sin\theta\\ -\sin\theta&\cos\theta\end{array}\right),$ (2)
@@
and controls the level of direct (diagonal) and cross-coupling (off-diagonal terms). We consider $\theta=\pi/2-0.1$, as values of $\theta$ close to $\pi/2$ have been shown to promote the existence of chimera states [22]. The small-world connectivity network is characterised by its $p-$dependant adjacency matrix $A^{p}$, whose elements satisfy $A_{ij}^{p}=\delta_{ij}k_{i}-\mathcal{L}_{ij}^{p}$, with $\delta_{ij}$ the Kroenecker delta, $k_{i}$ the degree (or number of connections) of the $i$th node, and $\mathcal{L}^{p}_{ij}$ the element $(i,j)$ of the Laplacian matrix of the network. Here we consider undirected small-world networks, so that the adjacency matrix is symmetric, and networks are constructed according to the Watts-Strogatz (WS) algorithm with rewiring probability $p$ and average degree $\langle k\rangle=6$ [33]. The results presented here remain mostly unchanged for different choices of $\langle k\rangle$.
I.2 Kuramoto order parameter
@@
Throughout this work, in order to characterise the synchronicity of the system, we compute the Kuramoto global order parameter [2]
@@
where $\theta_{j}(t)$ corresponds to a dynamical phase associated to the fast and slow variables $x_{j}(t)$ and $y_{j}(t)$ of the oscillator $j$, for $j=\left\{1,...,N\right\}$ defined as $\theta_{j}(t)~{}=~{}\arctan(y_{j}(t)/x_{j}(t))$. $R(t)\in[0,1]$, and gives relevant information about the collective dynamical regimes of the oscillators. Values of $R\approx 1$ correspond to highly synchronised states, with the limiting value $R=1$ accounting for global synchronisation, where $x_{i}=x_{j}$ and $y_{i}=y_{j}$ for all possible values of $i$ and $j$. When $R\ll 1$ the system is in an incoherent state, where oscillators do not share a common phase.
To study the phase space and the transitions between dynamical regimes as a function of the different parameters, we integrate Eq. (1) numerically (see Appendix A for details) and characterise the parameter space.
@@ II Results
Networks of FHN oscillators exhibit various dynamical regimes, including chaotic (Fig. 2a), intermittent (Fig. 2b), partially and fully synchronised (Fig. 2c and 2d, respectively) states, which depend not only of the strength of the coupling between oscillators ($d$) but also on the system size ($N$) and the network randomness, characterised by the rewiring probability ($p$). By increasing the coupling strength oscillators can alter their firing frequency, transiting from unimodal to bimodal distributions of inter-spike intervals (ISI), up to a threshold value where all oscillators become synchronised in their natural frequency (see right panels in Fig. 2). The bimodal ISI distributions (right panels in Fig. 2b and  2c) result from the existence of chimera states, where coherent and incoherent phases coexist.
II.1 Emergence of Epileptic-like events
It has been shown [11] that certain states exhibited by Eq. (1) resemble epileptic-like seizures in the brain, which are characterised by bursts of coherent activity separated by periods of incoherent activity (see Fig. 1a and 2b). Here we characterise the region of existence of these epileptic-like events for a range of network sizes, $N\in[10,250]$, rewiring probabilities, $p\in[0,1]$, and coupling strengths, $d\leq 0$ (Fig. 3).
Similarly to [11], we define an epileptic-like event as a state in which the order parameter $R(t)$ exceeds a threshold value $R_{\text{th}}=0.9$ for period of at leats $20$ simulation time units. With this, we measured the frequency of epileptic-like events (Ω=number of events/total $\Omega=\text{number of events}/\text{total time}$) as a function of the coupling strength and network size for different values of the network rewiring probability, $p$, which controls the randomness of the network (see results in Fig. 3a-c).
Our analyses reveal that network randomness is crucial for the emergence of epileptic-like events, which occur only within an intermediate range of coupling strengths (see yellow/red regions in Fig. 3b-c). Interestingly, these events are observed over a broader range of coupling strengths in larger networks, whereas smaller networks tend to exhibit more regular dynamics, limiting the occurrence of such intermittent behaviours. By definition, epileptic-like events correspond to chimera states, whose region of existence lies in the parameter space region limited by fully incoherent chaotic phase (low coupling strength) and a fully synchronised state (high coupling strength).
We note that the duration of these epileptic-like states deep in the intermittence region, followed a power-law like decay, despite of not being in a critical regime (see Fig. 3d). The time between epileptic-like events, on the other hand, followed an exponential decay, with a rate that depended on the parameters $d$ and $N$ (see Fig. 3e).

### v2

# Extreme events at the onset of epileptic-like intermittent activity of FitzHugh-Nagumo oscillators on small-world networks

@@ Abstract
In this work, we explore the influence of coupling strength, network size, and randomness on the collective dynamics of FitzHugh-Nagumo oscillators on complex networks. Using Watts-Strogatz small-world network connectivities, we identify four distinct dynamical phases: chaotic, intermittent, partially synchronized, and fully synchronized. The intermittent phase is characterized by the coexistence of chaotic behavior and chimera states, reminiscent of epileptic-seizure-related (ESR) intermittency observed in the brain. We analyze the inter-spike intervals of the individual oscillators, and the existence, duration, and frequency of ESR events as a function of the system parameters. Furthermore, we study the transitions into and out of the intermittent phase and show that peaks in the probability of extreme events—short transients of anomalously high synchronization—precede the transitions from chaos to intermittency and from partial to full synchronization. These transitions are followed by significant changes in the maximum Lyapunov exponent and Kaplan-Yorke dimension. Finally, we discuss how the coupling strength and network properties can be leveraged to control the system’s state and the potential applications of extreme event analysis in the study of neural data.
@@ I Introduction
Synchronization is a ubiquitous natural phenomenon, observed across diverse fields, ranging from physics and chemistry to biology, neuroscience, and even social systems [1, 2, 3, 4, 5]. In the brain, coordinated firing of distinct regions of the brain is crucial for effective information processing. However, when dysregulated, synchronized states can trigger epileptic seizures (see Fig. 1a), which disrupt normal brain function [6, 7]. Understanding the conditions that give rise to these pathologically coherent events, as well as the biological and physical mechanisms that regulate them, is therefore of considerable importance.
@@
The human brain is composed of billions of neurons, which are grouped and compartmentalized. These compartments form an intricate functional connectome. In the language of graphs, each region of the brain may be modeled as a single node, and connections between regions as edges that link pairs of nodes, such that the whole brain is reduced to a complex network [8, 9, 10, 11, 12] (see Fig. 1b). The brain connectivity is not homogeneous, and may present small-world organization [9, 5, 8, 13, 14], characterized by high connectivity and short paths between nodes.
A variety of mathematical models exist to describe the excitable behavior of neurons and brain regions. The paradigmatic FitzHugh-Nagumo (FHN) oscillator [15] offers a minimal framework for studying excitable systems. This two-variable model was originally introduced by FitzHugh as a modification of the Van der Pol oscillator and later demonstrated by Nagumo to be analogous to an electrical circuit [16, 17]. A single FHN oscillator may have excitable behavior or self-sustained oscillations. However, when coupled through a complex network topology, FHN oscillators may exhibit first or second-order phase transitions and a variety of dynamical behaviors, such as chaos and chimera states [18, 19, 20]. In these systems, the network topology, time delays [21, 22], the energy of the system [23, 24], and electromagnetic fluxes [25], among other factor, may alter the synchronization dynamics.
Chimera states, where coherent and incoherent states coexist, are of particular relevance for biological networks [26], as they provide a physical mechanism for partial synchronization, a feature of functioning neural circuits. Works on the FHN model have extensively studied the appearance of chimera states in non-locally coupled rings and complex networks, providing a deep understanding of the mechanisms and conditions under which chimera states emerge [27, 19, 28, 18, 29, 30, 31]. Additionally, there have been attempts to connect chimera states observed in FHN oscillators on networks to epileptic-seizures-related (ESR) synchronization in the brain [32, 11]. In these systems, the dynamic is intermittent, switching back and forth between chaotic and chimera states. Biologically, these chimeras correspond to periods of abnormal synchronous activity that result in a loss of normal brain function [11, 33, 34].
In an attempt to characterize the oscillatory behavior of theoretical and real neural systems, researchers have investigated the emergence of extreme events (EEs) in specific observables, such as the oscillation amplitude or level of synchronization [35, 36]. In this context, an EE occurs when the observable takes a value that is abnormally large compared to its typical behavior. Among other applications, EEs have been used to identify emergent phases in locally coupled Hindmarsh-Rose neurons [37] and in the bursting of neuronal activity [38, 39]. Thus, extreme event analysis provides a versatile tool to characterize complex system dynamics.
@@
Figure 1: Network representation of the brain. (a) Electroencephalogram (EEG) from a human brain exhibits synchronized regular activity between distinct regions of the brain, characteristic of an epileptic seizure event (adapted from [40]). (b) Anatomical regions of the brain (left). Regions are interpreted as nodes, and the connectivity network (right) is inferred from the neural activity (adapted from [14]). (c) WS networks for different values of the rewiring probability ($p$), color of the nodes indicate their degree ($k$). (d) Degree distribution for WS networks for different values of $p$ (increasing from left to right), and different network sizes (increasing from top to bottom).
Questions remain open regarding how real neural systems transit from a normal to an ESR intermittent state. Furthermore, understanding how this transition may be predicted and controlled may have a positive impact on the quality of life of people affected by related health conditions. In this work, we approach some of these questions from a theoretical perspective, using a minimal model of FHN oscillators on Watts-Strogatz (WS) complex networks [5]. These networks capture some of the small-world properties of the brain and allow for the investigation of the influence of network randomness through the rewiring probability parameter $p$, and network size on the collective dynamics [41] (see Fig. 1c and 1d).
@@
We show that the system of FHN oscillators on a WS network has four distinct dynamical phases: chaotic, intermittent, partially and fully synchronized. The intermittent phase exhibits transients of highly coherent, chimera states, separated in time by chaotic behavior, which were previously related to ESR events [11]. We find that the network size and the level of randomness of its connectivity, as well as the coupling strength between oscillators, control the emergence, duration and time between ESR events. Furthermore, we characterize the transition into and out of the zone of existence of chimeras. By analyzing the existence of EEs in the time series of the Kuramoto global order parameter, we show that the transitions from chaos to intermittency and the transition to full synchronization are preceded by a marked increase in the probability of observing short transients of anomalously high synchronization.
@@
The paper is organized as follows: In Secs. I.1 and II, we introduce the model and identify its dynamical regimes. In Sec. II.1, we show that the frequency and duration of ESR events depend on coupling strength, the number of oscillators, and network randomness. In Sec. II.2 we study the dynamical regimes and the stability of the synchronous solution. In Sec. II.3 show that the transitions between dynamical regimes can be characterized by the likelihood of observing EEs of synchronization. Finally, in Sec. III, we summarize and discuss our results and their potential relevance to biological systems.
@@
Figure 2: Dynamical regimes of FHN oscillators on WS networks. (a)-(d) Show the time series of the Kuramoto order parameter $R(t)$ (top), and the temporal evolution of the phases $\theta_{j}(t)\in[-\pi,\pi]$ for each of the $j=1,...,N=50$ oscillators (bottom), (e)-(h) show the distributions of inter-spike intervals (ISIs) for the $N=50$ oscillators, for couplings $d=0.01$ (a,e), $d=0.05$ (b,f), $d=0.10$ (c,g) and $d=0.22$ (d,h), in WS networks with $p=1$. In (a), the time window displayed is arbitrarily chosen to illustrate the dynamic regime of interest. An ESR event is highlighted in (b). In (f), red and blue curves show the ISI distribution for the chaotic regime and chimera transients, respectively.
I.1 FHN oscillators on complex networks
@@
The FitzHugh-Nagumo model comprises two variables: An activation variable, $x(t)$, capturing the rapid activation dynamics, and an inhibition variable, $y(t)$, representing the slower recovery process. In this study, we consider a network of $N$ coupled FHN oscillators, where the dynamics of each oscillator is governed by [42, 43, 30, 29]
$\left(\begin{array}[]{c}\varepsilon\dot{x}_{i}\\ \dot{y}_{i}\end{array}\right)=\left(\begin{array}[]{c}x_{i}-\frac{x_{i}^{3}}{3}-y_{i}\\ x_{i}+a\end{array}\right)+d\sum_{j=1}^{N}A^{p}_{ij}\bm{R}(\alpha)\left(\begin{array}[]{c}x_{j}-x_{i}\\ y_{j}-y_{i}\end{array}\right).$ (1)
@@
Here, $\dot{x}$ and $\dot{y}$ correspond to the time-derivatives of $x(t)$ and $y(t)$, respectively, and the time-scale separation $\varepsilon=0.05$ is fixed so that $x(t)$ has a fast dynamics compared to $y(t)$. The parameter $a$ controls the transition from excitable ($|a|>1$) to oscillatory ($|a|<1$) behavior. As our focus here is to study interactions of actively firing units, we set $a=0.5$. The summation term in Eq. (1) accounts for the inter-oscillator coupling, which is proportional to a uniform coupling strength $d$. The coupling matrix $\bm{R}(\alpha)$ is defined as
$\bm{R}(\alpha)=\left(\begin{array}[]{cc}\cos\alpha&\sin\alpha\\ -\sin\alpha&\cos\alpha\end{array}\right),$ (2)
@@
and controls the level of direct coupling (diagonal terms) and cross-coupling (off-diagonal terms). We consider $\alpha=\pi/2-0.1$, as values of $\alpha$ close to $\pi/2$ have been shown to promote the existence of chimera states [30]. The small-world network connectivity, for a given value of the rewiring probability $p$, is characterized by its adjacency matrix $A^{p}$, whose elements satisfy $A_{ij}^{p}=\delta_{ij}k_{i}-\mathcal{L}_{ij}^{p}$, with $\delta_{ij}$ the Kroenecker delta, $k_{i}$ the degree (or number of connections) of the $i$th node, and $\mathcal{L}^{p}_{ij}$ the element $(i,j)$ of the Laplacian matrix of the network. Here we consider undirected small-world networks, so that the adjacency matrix is symmetric, and networks are constructed according to the Watts-Strogatz (WS) algorithm with rewiring probability $p$ and average degree $\langle k\rangle=6$ [44] (see Fig. 1c and 1d). The results presented here remain largely unchanged for different choices of $\langle k\rangle$.
I.2 Kuramoto global order parameter
@@
Throughout this work, in order to characterize the synchronicity of the system, we compute the Kuramoto global order parameter [2]
@@
where $\theta_{j}(t)$ corresponds to a dynamical phase associated with the fast and slow variables $x_{j}(t)$ and $y_{j}(t)$ of the oscillator $j$ for $j=\left\{1,...,N\right\}$, defined as $\theta_{j}(t)~{}=~{}\arctan(y_{j}(t)/x_{j}(t))$. $R(t)\in[0,1]$, and gives relevant information about the collective dynamical regimes of the oscillators. Values of $R(t)=1$ account for global synchronization, where $x_{i}=x_{j}$ and $y_{i}=y_{j}$ for all nodes $i$ and $j$. Additionally, values $10<R(t)<$ 1 indicates states of incoherence and partial synchronization [45].
To study the phase space and the transitions between dynamical regimes as a function of the different parameters in the system, we integrate Eq. (1) numerically (see Appendix A for details) and characterize the $N-d-p$ parameter space.
@@ II Results
Networks of FHN oscillators exhibit various dynamical regimes, including chaotic (Fig. 2a), intermittent (Fig. 2b), partially and fully synchronized states (Fig. 2c and 2d, respectively), which depend not only on the strength of the coupling between oscillators ($d$) but also on the system size ($N$) and the network randomness (controlled by the rewiring probability $p$).
First, we measured the inter-spike-intervals (ISIs), defined as the time-intervals between consecutive peaks of the activation variable $x(t)$, for each oscillator. For fixed $N$ and $p$, a low (non-zero) coupling strength, $d$, results in incoherent dynamics, characterized by a unimodal distribution of ISIs (Fig. 2e). As the coupling strength is increased, the system enters the intermittent regime, where the dynamics of $R(t)$ switches between chaotic and chimera states. In this regime, the distribution of ISIs shifts to a bimodal behavior. Closer inspection of the distribution revealed that the first peak is associated with the chaotic phases (red curve in Fig. 2f), while the second peak corresponded to oscillations within the chimera states (blue curve in Fig. 2f). As the coupling is increased further, intermittency is lost and the system reaches partially and then fully synchronized states, both of which exhibit unimodal distributions of ISIs (Figs. 2g-h). In the partially synchronized regime (Figs. 2c and 2g), chimera states persist for the whole integration time, and their ISI is consistent with that the intermittent chimeras in Fig. 2f. In the fully synchronized phase all oscillators fire at their natural frequency, and the distribution of ISIs approximated a Dirac-delta function (Figs. 2d and 2h). We note that the dispersion in Fig. 2h comes from approximation errors in the dynamical phase analysis and localization fo the peaks.
II.1 Emergence of ESR synchronization events
It was previously shown [11] that transient chimera states observed in the intermittent regime of model (1) resemble ESR events of synchronization in the brain (see Fig. 1a and 2b). Here, we define an ESR event as a transient chimera state whose duration is of at least $20$ simulation time units. These states are detected by analyzing the envelope of the order parameter $R(t)$ and identifying all values that lie above a threshold $R_{\text{th}}=0.9$. This threshold corresponds approximately to the time-average of the order parameter $\bar{R}$ in the intermittent regime. We note that slight changes of $R_{\text{th}}$ do not affect our results. With this definition, we characterize the region of existence of ESR events for a range of network sizes, $N\in[10,250]$, rewiring probabilities, $p\in[0,1]$, and coupling strengths, $d\in[0,1]$ (Fig. 3).
[…]